OpenStudy (anonymous):

lim A_0 (1-cosA+sinA)/(1-cosA-sinA) if u r familiar and grown more up with limits please help me to get the answer

OpenStudy (sid1729):

As A approaches zero, both the numerator and denominator approach zero individually, so you can apply L'Hopital's rule here.

OpenStudy (anonymous):

can you please explain me the L'Hopital's rule if you are

OpenStudy (anonymous):

Look, you can treat numerator and denominator as two functions: \[1 - \cos a + \sin a\] and \[1 - \cos a - \sin a\] If you look here: http://en.wikipedia.org/wiki/L'H%C3%B4pital's_rule, you'll discover that we can compute limit by computing derivatives of numerator and denominator first.

OpenStudy (anonymous):

thanks alot